# Observation Interference in Partially Observable Assistance Games

### v1

@@ Abstract
We study partially observable assistance games (POAGs), a model of the human-AI value alignment problem which allows the human and the AI assistant to have partial observations. Motivated by concerns of AI deception, we study a qualitatively new phenomenon made possible by partial observability: would an AI assistant ever have an incentive to interfere with the human’s observations? First, we prove that sometimes an optimal assistant must take observation-interfering actions, even when the human is playing optimally, and even when there are otherwise-equivalent actions available that do not interfere with observations. Though this result seems to contradict the classic theorem from single-agent decision making that the value of perfect information is nonnegative, we resolve this seeming contradiction by developing a notion of interference defined on entire policies. This can be viewed as an extension of the classic result that the value of perfect information is nonnegative into the cooperative multiagent setting. Second, we prove that if the human is simply making decisions based on their immediate outcomes, the assistant might need to interfere with observations as a way to query the human’s preferences. We show that this incentive for interference goes away if the human is playing optimally, or if we introduce a communication channel for the human to communicate their preferences to the assistant. Third, we show that if the human acts according to the Boltzmann model of irrationality, this can create an incentive for the assistant to interfere with observations. Finally, we use an experimental model to analyze tradeoffs faced by the AI assistant in practice when considering whether or not to take observation-interfering actions.
@@ 1 Introduction
Assistance games provide a formalization of the human-AI value alignment problem [31]. They are based on Hidden Goal MDPs [12] and Cooperative Inverse Reinforcement Learning (CIRL) [15], an extension of Inverse Reinforcement Learning (IRL) [26, 1]. In assistance games, a single human and a single AI assistant share the same reward function, but this reward function is only known to the human; the assistant must learn it. In assistance games, desirable properties, such as teaching by the human and learning by the assistant, emerge as optimal solutions to the game [31]. (This contrasts with prior work on algorithms where teaching is an explicit objective [8, 14, 3].) For example, Woodward et al. [38] find that deep neural networks solving an assistance game invent strategies that involve information sharing, information seeking, and question answering.
Past analysis of assistance games was done assuming that the state of the world is fully observed by both the human and the assistant [15]. While Shah et al.’s [31] definition of an assistance game allows for partial observability, they do not study its implications. In this work, we introduce the notion of a partially observable assistance game (POAG) to study the more general case faced in reality: when the world is only partially observable. Partial observability raises new issues surrounding the communication of private information. A priori, we might hope that aligned AI assistants always give us complete information. Yet our analysis will show that even assistants which perfectly share our goals must make choices about what information to convey—and what information to obstruct.
This tension connects to broader work on AI deception, which recent research approaches from multiple angles. Park et al. [28] provide a philosophical definition and empirical survey of AI deception, while Ward et al. [33] define deception in structural causal games. Of particular relevance is work analyzing how reinforcement learning from human feedback (RLHF)—which can be seen as an algorithm for solving assistance games—can lead to deception. Lang et al. [22] prove that partial observability in RLHF can create dual risks of deceptive inflating and overjustification. Complementing Lang et al. [22]’s theory, Wen et al. [35], Williams et al. [37] provide experimental evidence that optimizing for human feedback teaches language models to mislead humans. However, these works primarily focus on misaligned AI systems that deceive for their own goals. We study the subtle case where a perfectly aligned AI assistant might obstruct information for the human’s benefit.
@@
Our analysis reveals three distinct incentives for an AI assistant to take observation interfering actions. First, when the assistant has private information, it might need to interfere with observations to communicate its private information to the human (Section 4.2). This can happen even when the human is playing optimally, and even when there are otherwise-equivalent actions available that do not interfere with observations. This result presents a puzzle, as it seems to contradict the classic theorem from single-agent decision making that the value of perfect information is nonnegative. To resolve this seeming contradiction, we develop a notion of interference defined on entire policies rather than individual actions. While optimal solutions (i.e., human-AI policy pairs) might involve the AI assistant taking individual actions which would on their own constitute observation interference, we prove that there is always an optimal solution with no observation interference when we consider the AI assistant’s overall policy. This can be viewed as an extension of the classic result that the value of perfect information is nonnegative into the cooperative multiagent setting.
This result connects to a broader literature on the value of information in multiagent settings. In games with competing interests, it is well-known that introducing common knowledge can lead to worse outcomes for all players [19]. Using a set-theoretic framework, Bassan et al. [4] establish a class of general-sum games where additional information Pareto-improves all of the Nash equilibria. Their class of games includes common-payoff games. Using a probabilistic framework, Lehrer et al. [23] extend this analysis to alternative solution concepts. Notably, Bassan et al. [4], Lehrer et al. [23] consider only single-timestep games where players simultaneously act without observing the other players’ actions. In our setting, the environment evolves over time, and the players can observe each other’s actions to make better inferences about the state of the world. Our results show that observing the actions of other players is a key feature that enables observation interference to communicate private information and achieve better outcomes.
@@
Finally, in Section 7, we use an experimental model to investigate tradeoffs the assistant must make when considering whether or not to interfere with observations. In line with our theory, we find that observation interference allows the AI assistant to communicate private information, but it comes at the cost of destroying useful information. Measuring this tradeoff, we find that having more private information leads to a stronger incentive to interfere with observations.
Our results establish that optimal assistants might need to interfere with observations in optimal policy pairs as well as when responding optimally to fixed human policies. By the definition of optimality, all the cases of observation interference that we identify are beneficial to the human. In practice, however, the assistant might be imperfectly aligned, and it might be acting suboptimally. In these cases, observation interference might be detrimental to the human. We intend for our theoretical characterization of interference in optimal solutions to establish a framework that can help distinguish between different forms of observation interference in practice.
@@ 2.1 Partially Observable Assistance Games
We study partially observable assistance games [31]:
@@
A partially observable assistance game (POAG) $M$ is a two-player DecPOMDP with a human or principal, $\mathbf{H}$, and an AI assistant, $\mathbf{A}$. The game is described by a tuple, $M~{}=~{}\langle\mathcal{S},\{\mathcal{A}^{\mathbf{H}},\mathcal{A}^{\mathbf{A}}\},T(\cdot\mid\cdot,\cdot,\cdot),\{\Theta,R(\cdot,\cdot,\cdot;\cdot)\},\{{\Omega}^{\mathbf{H}},{\Omega}^{\mathbf{A}}\},O(\cdot,\cdot\mid\cdot,\cdot,\cdot),P_{0}(\cdot,\cdot),\gamma\rangle,$ with the following definitions:
  • $\mathcal{S}$
    a set of world states: $s\in\mathcal{S}$.
  • $\mathcal{A}^{\mathbf{H}}$
    a set of actions for $\mathbf{H}$: $a^{\mathbf{H}}\in\mathcal{A}^{\mathbf{H}}$.
  • $\mathcal{A}^{\mathbf{A}}$
    a set of actions for $\mathbf{A}$: $a^{\mathbf{A}}\in\mathcal{A}^{\mathbf{A}}$.
  • $T(\cdot\mid\cdot,\cdot,\cdot)$
    a conditional distribution on the next world state, given previous state and action for both players: $T(s^{\prime}\mid s,a^{\mathbf{H}},a^{\mathbf{A}})$.
  • $\Theta$
    a set of possible static reward parameter values, only observed by $\mathbf{H}$: $\theta\in\Theta$.
  • $R(\cdot,\cdot,\cdot;\cdot)$
    a parameterized reward function that maps world states, joint actions, and reward parameters to real numbers. $R:\mathcal{S}\times\mathcal{A}^{\mathbf{H}}\times\mathcal{A}^{\mathbf{A}}\times\Theta\rightarrow\mathbb{R}.$
  • ${\Omega}^{\mathbf{H}}$
    a set of observations for $\mathbf{H}$: $o^{\mathbf{H}}\in{\Omega}^{\mathbf{H}}$.
  • ${\Omega}^{\mathbf{A}}$
    a set of observations for $\mathbf{A}$: $o^{\mathbf{A}}\in{\Omega}^{\mathbf{A}}$.
  • $O(\cdot,\cdot\mid\cdot,\cdot,\cdot)$
    a conditional distribution on the observations, given the next world state and action of both players: $O(o^{\mathbf{H}},o^{\mathbf{A}}\mid s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}})$.
  • $P_{0}(\cdot,\cdot)$
    a distribution over the initial state, represented as tuples: $P_{0}(s_{0},\theta)$.
  • $\gamma$
    a discount factor: $\gamma\in[0,1]$.
@@
Note that optimal policy pairs are in particular Nash equilibria. Computationally, POAGs are equivalent to 2-player DecPOMDPs. Thus, finding optimal policy pairs for POAGs is NEXP-hard in general [5] [cf. 30]. A POAG may have multiple distinct optimal policy pairs. For instance, $\mathbf{H}$ and $\mathbf{A}$ may have multiple ways of resolving coordination problems, or choose different ways of communicating.
@@ 2.2 Beliefs and calibration of beliefs
To address this question, we apply known techniques to establish what information $\mathbf{H}$ needs to form calibrated beliefs in a POAG. The key idea is that if $\mathbf{H}$ knows $\mathbf{A}$’s policy, $\mathbf{H}$ can treat $\mathbf{A}$ like any other part of the environment. Forming beliefs then reduces to POMDP inference.
The simplest case of $\mathbf{H}$ knowing $\mathbf{A}$’s policy is when $\mathbf{A}$ is playing a fixed policy:
@@
In Section 4, we study when observation interference occurs in optimal policy pairs, i.e., when $\mathbf{H}$ and $\mathbf{A}$ are each playing a best response to the other. By design, this solution concept assumes that $\mathbf{H}$ knows whatever information is needed about $\mathbf{A}$’s policy to compute a best response. In such a case where $\mathbf{H}$ knows $\mathbf{A}$’s policy, the preceding results show that $\mathbf{H}$ can form calibrated beliefs about the world, even when $\mathbf{A}$ is interfering with observations. Observation interference increases $\mathbf{H}$’s uncertainty, but it doesn’t break the calibration of $\mathbf{H}$’s beliefs. Because $\mathbf{H}$ can still form calibrated beliefs in this setting, our work uses the concept of “interference” rather than the concept of “deception.”
@@ 3 Defining observation interference
First, we define what interference means. Intuitively, interference is taking action so that the human receives a less informative signal about the state. In particular, the human receives, in some sense, a subset of the information. We formalize this by saying one signal is less informative than another about the state if (without knowing the state) we could generate one signal from the other [cf. 7, 6, 10].
@@
Let $(P(\cdot\mid s))_{s\in\mathcal{S}}$ and $(\hat{P}(\cdot\mid s))_{s\in\mathcal{S}}$ be families of probability distributions over $\Omega$. We say that $\hat{P}$ is at most as informative as $P$ if there exists a stochastic function $F\colon\Omega\rightsquigarrow\Omega$ s.t. for all states $s$ we have $F(X)\sim\hat{P}(\cdot\mid s)$ if $X\sim P(\cdot\mid s)$. We say that $P$ is (strictly) more informative than $\hat{P}$ if $P$ is at least as informative as $\hat{P}$ but not vice versa.
Why do we include the condition “for all states $s$” in Definition 3.1? Intuitively, we want it always to be possible to use the stochastic function $F$ to reconstruct the less informative signal from the more informative signal. Since our setting is partially observable, the “for all states $s$” condition allows a player of the game to do this reconstruction in any scenario, even if their observations don’t enable them to infer the state.
Note that this definition induces only a partial order on probability distributions. For instance, different signals may provide information about different aspects of $s$, and it may not be possible to generate either distribution from the other.
With this definition in hand, we define an observation-interfering action as one that results in the human’s observation being less informative about the state than the observation distribution resulting from another assistant action. We additionally require that this other action has the same effects on the state and immediate reward. After all, it is clear that sometimes $\mathbf{A}$ has to trade off providing information to $\mathbf{H}$ with optimizing its effect on the environment. Formally:
@@
The definition may be refined in various ways. For instance, note that the above does not take into account the information that the human has other than the current $o^{\mathbf{H}}$. Arguably, removing a signal that $\mathbf{H}$ can reconstruct from her past observations should not be viewed as signal interference. Our definition does not align with this judgment. However, none of these modifications matter for our analysis below. Thus, we have opted for the simplest definition. We discuss these in more detail in Appendix G.
@@
To understand the conditions under which interference occurs, it is useful to consider POAGs in which one of the players has no private information.
@@
For a POAG $M$, we say $\mathbf{A}$ has no private information if there exists a function $f$ determining $\mathbf{A}$’s observations from $\mathbf{H}$’s observations. For all state-action tuples $(s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}})$ and observation pairs $(o^{\mathbf{H}},o^{\mathbf{A}})\in\emph{\text{supp}}(O(\cdot,\cdot\mid s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}}))$, then $f$ must satisfy $f(o^{\mathbf{H}})=o^{\mathbf{A}}$.
@@
To construct $M^{\mathbf{H}\rightarrow\mathbf{A}}$, let $\mathcal{M}$ be some set of possible messages/signals s.t. there is an injection ${\Omega}^{\mathbf{A}}\xhookrightarrow{}\mathcal{M}$. Then, construct a new human action space ^$\hat{}\mathcal{A}^{\mathbf{H}}=\mathcal{A}^{\mathbf{H}}\times\mathcal{M}$ and new assistant observation space ^$\hat{}{\Omega}^{\mathbf{A}}={\Omega}^{\mathbf{A}}\times\mathcal{M}$. The new observation kernel has $\hat{O}\left(o^{\mathbf{H}},(o^{\mathbf{A}},m^{\prime})\mid s^{\prime},(a^{\mathbf{H}},m),a^{\mathbf{A}}\right)=\mathbb{1}[m{=}m^{\prime}]O(o^{\mathbf{H}},o^{\mathbf{A}}\mid s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}})$. For everything else, the messages are simply ignored.
@@
We may have various expectations on how $\mathbf{H}$ will play in a POAG. Especially if there are multiple optimal policy pairs, we may expect some of these policy pairs to be more plausible because they require simpler behavior of the human [cf. 17, 32]. Both of the conditions below are based on the idea that $\mathbf{A}$ and $\mathbf{H}$ are unlikely to use consequential actions in the world to communicate with each other.
Our first condition intends to express a form of naivete on $\mathbf{H}$’s part in how she interprets her observations. Roughly, the condition says that $\mathbf{H}$ takes her observations at face value, i.e., as if they were not interfered with. She does not try to interpret them as a form of communication by $\mathbf{A}$. For instance, if $\mathbf{H}$ reads a thermometer as saying that a temperature is $37$ degrees, she chooses under the assumption that the temperature is indeed $37$ degrees, rather than, say, interpreting $37$ as a message sent by $\mathbf{A}$ which may have interfered with the thermometer.
@@
The second property is that when the human knows that her action has no effect on the state, then she chooses among actions that maximize immediate reward. To state this formally, we first define the following. We say that in $h_{t}^{\mathbf{H}}$ actions don’t affect state transitions, if for all $s$ s.t. we have $P(s\mid h_{t}^{H},\pi^{\mathbf{A}})>0$ for some $\pi^{\mathbf{A}}$, we have that for all $a^{\mathbf{A}}$ $P(s^{\prime}\mid s,a^{\mathbf{A}},a^{\mathbf{H}})$ is constant over $a^{\mathbf{H}}$. We say that $\pi^{\mathbf{H}}$ myopically maximizes reward in $h_{t}^{\mathbf{H}}$ if we have that there is some distribution $\alpha^{\mathbf{A}}\in\Delta(\mathcal{A}^{\mathbf{A}})$ s.t. $\pi^{\mathbf{H}}(\cdot\mid h)$ randomizes only over actions in $\operatorname*{arg\,max}_{a^{\mathbf{H}}}\mathbb{E}_{a^{\mathbf{A}}\sim\alpha^{\mathbf{A}},s\sim P(\cdot\mid h,a^{\mathbf{H}},a^{\mathbf{A}})}\left[R(s,a^{\mathbf{H}},a^{\mathbf{A}},\theta)\right]$.
@@
We say that a human policy $\pi^{\mathbf{H}}$ acts naively if whenever $\mathbf{H}$ faces a choice that doesn’t affect state transitions (but potentially an effect on $\mathbf{A}$’s observation), $\mathbf{H}$ plays an action that myopically maximizes reward.
Intuitively, $\alpha^{\mathbf{A}}$ is $\mathbf{H}$’s belief about what action $\mathbf{A}$ is going to take. Importantly, if $\mathbf{H}$ acts naively, she is unwilling to play a suboptimal action in order to communicate information to $\mathbf{A}$.

### v2

@@ Abstract
We study partially observable assistance games (POAGs), a model of the human-AI value alignment problem which allows the human and the AI assistant to have partial observations. Motivated by concerns of AI deception, we study a qualitatively new phenomenon made possible by partial observability: would an AI assistant ever have an incentive to interfere with the human’s observations? First, we prove that sometimes an optimal assistant must take observation-interfering actions, even when the human is playing optimally, and even when there are otherwise-equivalent actions available that do not interfere with observations. Though this result seems to contradict the classic theorem from single-agent decision making that the value of information is nonnegative, we resolve this seeming contradiction by developing a notion of interference defined on entire policies. This can be viewed as an extension of the classic result that the value of information is nonnegative into the cooperative multiagent setting. Second, we prove that if the human is simply making decisions based on their immediate outcomes, the assistant might need to interfere with observations as a way to query the human’s preferences. We show that this incentive for interference goes away if the human is playing optimally, or if we introduce a communication channel for the human to communicate their preferences to the assistant. Third, we show that if the human acts according to the Boltzmann model of irrationality, this can create an incentive for the assistant to interfere with observations. Finally, we use an experimental model to analyze tradeoffs faced by the AI assistant in practice when considering whether or not to take observation-interfering actions.
assistance games, AI alignment, observation interference, partial observability
@@ 1 Introduction
Assistance games provide a formalization of the human-AI value alignment problem (Shah et al., 2020). They are based on Hidden Goal MDPs (Fern et al., 2014) and Cooperative Inverse Reinforcement Learning (CIRL) (Hadfield-Menell et al., 2016), an extension of Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000; Abbeel & Ng, 2004). In assistance games, a single human and a single AI assistant share the same reward function, but this reward function is only known to the human; the assistant must learn it. In assistance games, desirable properties, such as teaching by the human and learning by the assistant, emerge as optimal solutions to the game (Shah et al., 2020). (This contrasts with prior work on algorithms where teaching is an explicit objective (Cakmak & Lopes, 2012; Goldman & Kearns, 1995; Balbach & Zeugmann, 2009).) For example, Woodward et al. (2020) find that deep neural networks solving an assistance game invent strategies that involve information sharing, information seeking, and question answering.
Past analysis of assistance games was done assuming that the state of the world is fully observed by both the human and the assistant (Hadfield-Menell et al., 2016, 2017). While Shah et al.’s (2020) definition of an assistance game allows for partial observability, they do not study its implications. In this work, we introduce the notion of a partially observable assistance game (POAG) to study the more general case faced in reality: when the world is only partially observable. Partial observability raises new issues surrounding the communication of private information. A priori, we might hope that AI assistants never take any action that obstructs information. Yet our analysis will show that even assistants which perfectly share our goals must sometimes obstruct information to communicate other, more important information.
This tension connects to broader work on AI deception, which recent research approaches from multiple angles. Park et al. (2024) provide a philosophical definition and empirical survey of AI deception, while Ward et al. (2023) define deception in structural causal games. Of particular relevance is work analyzing how reinforcement learning from human feedback (RLHF)—which can be seen as an algorithm for solving assistance games—can lead to deception. Lang et al. (2024) prove that partial observability in RLHF can create dual risks of deceptive inflating and overjustification. Complementing Lang et al. (2024)’s theory, Wen et al. (2024) and Williams et al. (2024) provide experimental evidence that optimizing for human feedback teaches language models to mislead humans. However, these works primarily focus on misaligned AI systems that deceive for their own goals. We study the subtle case where a perfectly aligned AI assistant might obstruct information for the human’s benefit.
@@
Our analysis reveals three distinct incentives for an AI assistant to take observation interfering actions. First, when the assistant has private information, it might need to interfere with observations to communicate its private information to the human (Section 4.2). This can happen even when the human is playing optimally, and even when there are otherwise-equivalent actions available that do not interfere with observations. This result presents a puzzle, as it seems to contradict the classic theorem from single-agent decision making that the value of information (sometimes also called the value of perfect information) is nonnegative (e.g., Koller & Friedman, 2009, Sect. 23.7; Russell & Norvig, 2010, Sect. 16.6.3). To resolve this seeming contradiction, we develop a notion of interference defined on entire policies rather than individual actions. While optimal solutions (i.e., human-AI policy pairs) might involve the AI assistant taking individual actions which would on their own be observation interference, we prove that there is always an optimal solution with no observation interference when we consider the AI assistant’s overall policy. This can be viewed as an extension of the classic result that the value of (perfect) information is nonnegative into the cooperative multiagent setting.
This result connects to a broader literature on the value of information in multiagent settings. In games with competing interests, it is well-known that introducing common knowledge can lead to worse outcomes for all players (Kamien et al., 1990). Using a set-theoretic framework, Bassan et al. (2003) establish a class of general-sum games where additional information Pareto-improves all of the Nash equilibria. Their class of games includes common-payoff games. Using a probabilistic framework, Lehrer et al. (2010) extend this analysis to alternative solution concepts. Notably, Bassan et al. (2003) and Lehrer et al. (2010) consider only single-timestep games where players simultaneously act without observing the other players’ actions. In our setting, the environment evolves over time, and the players can influence each other’s observations through their actions. Our results show that this influence on observations—including observation interference—is a key feature that enables the communication of private information to achieve better outcomes.
@@
Finally, in Section 7, we use an experimental model to investigate tradeoffs the assistant faces when deciding whether or not to interfere with observations. In line with our theory, we find that observation interference allows the AI assistant to communicate private information, but it comes at the cost of destroying useful information. Measuring this tradeoff, we find that having more private information leads to a stronger incentive to interfere with observations.
@@ 2.1 Partially Observable Assistance Games
We study partially observable assistance games (Shah et al., 2020):
@@
A partially observable assistance game (POAG) $M$ is a two-player DecPOMDP with a human or principal, $\mathbf{H}$, and an AI assistant, $\mathbf{A}$. The game is described by a tuple, $M~=~\langle\mathcal{S},\{\mathcal{A}^{\mathbf{H}},\mathcal{A}^{\mathbf{A}}\},T(\cdot\mid\cdot,\cdot,\cdot),\{\Theta,R(\cdot,\cdot,\cdot;\cdot)\},\{{\Omega}^{\mathbf{H}},{\Omega}^{\mathbf{A}}\},\text{$O(\cdot,\cdot\mid\cdot,\cdot,\cdot)$},P_{0}(\cdot,\cdot),\gamma\\ \rangle,$ with the following definitions: $\mathcal{S}$, a set of world states: $s\in\mathcal{S}$; $\mathcal{A}^{\mathbf{H}}$, a set of actions for $\mathbf{H}$: $a^{\mathbf{H}}\in\mathcal{A}^{\mathbf{H}}$; $\mathcal{A}^{\mathbf{A}}$, a set of actions for $\mathbf{A}$: $a^{\mathbf{A}}\in\mathcal{A}^{\mathbf{A}}$; $T(\cdot\mid\cdot,\cdot,\cdot)$, a conditional distribution on the next world state, given previous state and action for both players: $T(s^{\prime}\mid s,a^{\mathbf{H}},a^{\mathbf{A}})$; $\Theta$, a set of possible static reward parameter values, only observed by $\mathbf{H}$: $\theta\in\Theta$; $R(\cdot,\cdot,\cdot;\cdot)$, a parameterized reward function that maps world states, joint actions, and reward parameters to real numbers: $R:\mathcal{S}\times\mathcal{A}^{\mathbf{H}}\times\mathcal{A}^{\mathbf{A}}\times\Theta\rightarrow\mathbb{R}$; ${\Omega}^{\mathbf{H}}$, a set of observations for $\mathbf{H}$: $o^{\mathbf{H}}\in{\Omega}^{\mathbf{H}}$; ${\Omega}^{\mathbf{A}}$, a set of observations for $\mathbf{A}$: $o^{\mathbf{A}}\in{\Omega}^{\mathbf{A}}$; $O(\cdot,\cdot\mid\cdot,\cdot,\cdot)$, a conditional distribution on the observations, given the next world state and action of both players: $O(o^{\mathbf{H}},o^{\mathbf{A}}\mid s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}})$; $P_{0}(\cdot,\cdot)$, a distribution over the initial state, represented as tuples: $P_{0}(s_{0},\theta)$; and $\gamma$, a discount factor: $\gamma\in[0,1]$.
@@
Note that optimal policy pairs are in particular Nash equilibria for the shared reward function $R$. Computationally, POAGs are equivalent to 2-player decentralized partially observable Markov decision processes (DecPOMDPs). Thus, finding optimal policy pairs for POAGs is NEXP-hard in general (Bernstein et al., 2002) (cf. Reif, 1984). A POAG may have multiple distinct optimal policy pairs, as there may be different ways for $\mathbf{H}$ and $\mathbf{A}$ to communicate or resolve coordination problems.
While the examples in this paper are simple, POAGs—and thus all our positive results—inherit the broad generality of DecPOMDPs. POAGs can model games where $\mathbf{H}$ acts first, where $\mathbf{A}$ acts first, or where $\mathbf{H}$ and $\mathbf{A}$ act simultaneously. POAGs allow both $\mathbf{H}$ and $\mathbf{A}$ to observe private information at multiple times, as well as take actions that influence both the state of the world and each other’s observations.
@@ 2.2 Beliefs and Calibration of Beliefs
To address this question, we apply known techniques to establish what information $\mathbf{H}$ needs to form calibrated beliefs in a POAG. (See Appendix A for proofs.) The simplest case of $\mathbf{H}$ knowing $\mathbf{A}$’s policy is when $\mathbf{A}$ is playing a fixed policy:
@@
When $\mathbf{H}$ knows $\mathbf{A}$’s policy, the preceding results show that $\mathbf{H}$ can form calibrated beliefs about the world, even when $\mathbf{A}$ is interfering with observations. Observation interference increases $\mathbf{H}$’s uncertainty, but it doesn’t break the calibration of $\mathbf{H}$’s beliefs. Because $\mathbf{H}$ can still form calibrated beliefs in this setting, our work uses the concept of “interference” rather than the concept of “deception.”
@@ 3 Defining Observation Interference
First, we define what interference means. Intuitively, interference is taking action so that the human receives a less informative signal about the state. In particular, the human receives, in some sense, a subset of the information. We formalize this by saying one signal is less informative than another about the state if (without knowing the state) we could generate one signal from the other (cf. Blackwell et al., 1951; Blackwell, 1953; de Oliveira, 2018).
@@
Let $(P(\cdot\mid s))_{s\in\mathcal{S}}$ and $(\hat{P}(\cdot\mid s))_{s\in\mathcal{S}}$ be families of probability distributions over $\Omega$. We say that $\hat{P}$ is at most as informative as $P$ if there exists a stochastic function $F\colon\Omega\rightsquigarrow\Omega$ (mapping observations to random variables over observations) s.t. for all states $s$ we have $F(X)\sim\hat{P}(\cdot\mid s)$ if $X\sim P(\cdot\mid s)$. We say that $P$ is (strictly) more informative than $\hat{P}$ if $P$ is at least as informative as $\hat{P}$ but not vice versa.
Why do we include the condition “for all states ss” in Definition 3.1? Intuitively, we want it always to be possible to use the stochastic function $F$ to reconstruct the less informative signal from the more informative signal. Since our setting is partially observable, the “for all states ss” condition allows a player to do this reconstruction in any scenario, even if their observations don’t enable them to infer the state.
With this definition in hand, we define an observation-interfering action as one that results in the human’s observation being less informative about the state than the observation distribution resulting from another assistant action. We additionally require that this other action has the same effects on the state and immediate reward. After all, it is clear that sometimes $\mathbf{A}$ has to trade off providing information to $\mathbf{H}$ with optimizing its effect on the environment.
@@
To understand the conditions under which interference occurs, it is useful to consider POAGs where one of the players has no private information.
@@
For a POAG $M$, we say $\mathbf{A}$ has no private information if there exists a function $f$ determining $\mathbf{A}$’s observations from $\mathbf{H}$’s observations. For all state-action tuples $(s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}})$ and observation pairs $(o^{\mathbf{H}},o^{\mathbf{A}})\in\emph{\text{supp}}(O(\cdot,\cdot\mid s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}}))$, then $f$ must have $f(o^{\mathbf{H}})=o^{\mathbf{A}}$.
@@
To construct $M^{\mathbf{H}\rightarrow\mathbf{A}}$, let $\mathcal{M}$ be some set of possible messages/signals s.t. there is an injection ${\Omega}^{\mathbf{A}}\xhookrightarrow{}\mathcal{M}$. Then, construct a new human action space $\hat{}\mathcal{A}^{\mathbf{H}}=\mathcal{A}^{\mathbf{H}}\times\mathcal{M}$ and new assistant observation space $\hat{}{\Omega}^{\mathbf{A}}={\Omega}^{\mathbf{A}}\times\mathcal{M}$. The new observation kernel has $\hat{O}\left(o^{\mathbf{H}},(o^{\mathbf{A}},m^{\prime})\mid s^{\prime},(a^{\mathbf{H}},m),a^{\mathbf{A}}\right)=\mathbb{1}[m{=}m^{\prime}]O(o^{\mathbf{H}},o^{\mathbf{A}}\mid s^{\prime},a^{\mathbf{H}},a^{\mathbf{A}})$. For everything else, the messages are simply ignored.
@@
We may have various expectations on how $\mathbf{H}$ will play in a POAG. Especially if there are multiple optimal policy pairs, we may expect some of these policy pairs to be more plausible because they require simpler behavior of the human (cf. Hu et al., 2020; Treutlein et al., 2021). Both of the conditions below are based on the idea that $\mathbf{A}$ and $\mathbf{H}$ are unlikely to use consequential actions in the world to communicate with each other.
Our first condition intends to express a form of naivete on $\mathbf{H}$’s part in how she interprets her observations. Roughly, the condition says that $\mathbf{H}$ takes her observations at face value, i.e., as if they were not interfered with. She does not try to interpret them as a form of communication by $\mathbf{A}$. For instance, if $\mathbf{H}$ reads a thermometer as saying that a temperature is $37$ degrees, she chooses under the assumption that the temperature is indeed $37$ degrees, rather than, say, interpreting $37$ as a message sent by $\mathbf{A}$.
@@
The second property is that when the human knows that her action has no effect on the state, then she chooses among actions that maximize immediate reward. To state this formally, we first define the following. We say that in $h_{t}^{\mathbf{H}}$ actions don’t affect state transitions, if for all $s$ s.t. we have $P(s\mid h_{t}^{H},\pi^{\mathbf{A}})>0$ for some $\pi^{\mathbf{A}}$, we have that for all $a^{\mathbf{A}}$ the transition probability $P(s^{\prime}\mid s,a^{\mathbf{A}},a^{\mathbf{H}})$ is constant over $a^{\mathbf{H}}$. We say that $\pi^{\mathbf{H}}$ myopically maximizes reward in $h_{t}^{\mathbf{H}}$ if there is some distribution $\alpha^{\mathbf{A}}\in\Delta(\mathcal{A}^{\mathbf{A}})$ s.t. $\pi^{\mathbf{H}}(\cdot\mid h_{t}^{\mathbf{H}})$ randomizes only over actions in $\operatorname*{arg\,max}_{a^{\mathbf{H}}}\mathbb{E}_{a^{\mathbf{A}}\sim\alpha^{\mathbf{A}},s\sim P(\cdot\mid h_{t}^{\mathbf{H}},a^{\mathbf{H}},a^{\mathbf{A}})}\left[R(s,a^{\mathbf{H}},a^{\mathbf{A}},\theta)\right]$. (Intuitively, $\alpha^{\mathbf{A}}$ is $\mathbf{H}$’s belief about what action $\mathbf{A}$ is going to take.)
@@
We say that a human policy $\pi^{\mathbf{H}}$ acts naively if whenever $\mathbf{H}$ faces a choice that doesn’t affect state transitions (but potentially affects $\mathbf{A}$’s observation), $\mathbf{H}$ plays an action that myopically maximizes reward.
Importantly, if $\mathbf{H}$ acts naively, she is unwilling to play a suboptimal action to communicate information to $\mathbf{A}$.